The sum of my mother’s and my father’s ages is 70 and my father is two years older than my mother. With my age added, our total is 80. How old will we all be when our total age is 101?

Answer by ankor@dixie-net.com(22740)

let a = my age
let m = Mom's age
let f = Dad's age
:
write an equation for each statement
:
The sum of my mother’s and my father’s ages is 70 and
m + f = 70
:
my father is two years older than my mother.
f = m + 2
:
With my age added, our total is 80.
a + m + f = 80
replace m+f with 70
a + 70 = 80
a = 10 yrs is your age
:
In the first equation, m + f = 70,replace f with m+2
m + 2 + m = 70
2m = 70 - 2
2m = 68
m = 68/2
m = 34 is Mom's age
then obviously
f = 34 + 2
f = 36 is Dad's age
:
How old will we all be when our total age is 101?
let y = no. of years for this to be true
(a+y) + (m+y) + (f+y) = 101
a + m + f + 3y = 101
Replace a+m+f with 80
80 + 3y = 101
3y = 101 - 80
3y = 21
y = 21/3
y = 7 yrs
Add 7 the each of their ages and you have it, then check the total to see if it adds up to 101